OpenStudy (anonymous):

If X varies directly as Y, and X=30 when Y=5, find X when Y=8.

OpenStudy (anonymous):

The last person who answered this for me was wrong

OpenStudy (anonymous):

X/Y = 30/5 = ?/8 So 48

OpenStudy (anonymous):

any graph or picture?

OpenStudy (anonymous):

THANK YOU SIR